OpenStudy (anonymous):

For questions 11 – 14, find the value of x in the given ratio.

OpenStudy (anonymous):



OpenStudy (anonymous):

Oh god you're going to draw; heh \[\left(\begin{matrix}2 \\3\end{matrix}\right)=\left(\begin{matrix}x \\ 15\end{matrix}\right)\] We are attempting to isolate x; so we multiple out the denominator \[15*\left(\begin{matrix}2 \\ 3\end{matrix}\right)=x\] 15*2 = 30; 30/3 = 10 10=x

OpenStudy (anonymous):

lol thanks for the drawing compliment -.-

OpenStudy (anonymous):

jk ok anyways haha